# Exploring Parent’s Needs for Children-Centered AI to Support Preschoolers’ Storytelling and Reading Activities

### v1

@@ Abstract.
Interactive storytelling is vital for preschooler development. While children’s interactive partners have traditionally been their parents and teachers, recent advances in artificial intelligence (AI) have sparked a surge of AI-based storytelling technologies. As these technologies become increasingly ubiquitous in preschoolers’ lives, questions arise regarding how they function in practical storytelling scenarios and, in particular, how parents, the most critical stakeholders, experience and perceive these technologies. This paper investigates these questions through a qualitative study with 17 parents of children aged 3-6. Our findings suggest that even though AI-based storytelling technologies provide more immersive and engaging interaction, they still cannot meet parents’ expectations due to a series of interactive, functional, and algorithmic challenges. We elaborate on these challenges and discuss the possible implications of future AI-based storytelling technologies for preschoolers. We conclude by highlighting the design implications for future AI-based storytelling technologies.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June XX–$X$, 20XX; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Interactive storytelling together with the parents is a critical step in the development for preschoolers111”Preschoolers” in this paper refers to young children, typically aged 3 to 6, who are in the stage of early childhood education that precedes formal schooling (of Disease Control and Prevention, 2021). (Glassner, 2017). Because preschoolers are often not able to decode text independently, parents often sit together with and read to them, ideally engaging the children in guided conversation (Zevenbergen and Whitehurst, 2003) by asking questions and providing responsive feedback. This process has been proven to be beneficial to children’s later language and literacy proficiency, communication skills, cultural and emotional awareness, and other aspects of cognitive development (Wright, 1995; Yen et al., 2018; Chang and Huang, 2016; Peck, 1989). However, due to various practical challenges, today’s parents may not always have sufficient language skills, knowledge, time, or inclination to engage in such conversation-rich storybook reading with their preschoolers, even though they believe in the importance of interactive storytelling with their children (Manz et al., 2010; Xu et al., 2021b).
In recent years, AI-based applications with multiple functions, materials, modalities, and interaction techniques have been becoming increasingly popular in supporting storytelling practices for preschoolers in the household setting. These applications could be embedded as applications in mobile phones or tablets (Dietz et al., 2021; Zhang et al., 2022), or exist with physical tangible presences (e.g., Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021)); they rely on interactive technologies tailored to children (e.g., (Xu et al., 2022; Yao et al., 2022)), or general voice assistant technologies (e.g., Amazon Alexa, Google Assistant, Apple Siri); they provide specialized features designed specifically for children’s storytelling (e.g., (Dietz et al., 2021; Zhang et al., 2022; Michaelis and Mutlu, 2017)), or carry out children’s instructions based on the embedded general voice assistant technologies (e.g., Amazon Alexa, Google Assistant, Apple Siri); some of these applications are mature commercial products (e.g., Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021)) and some others are still in the research period (e.g., storycoder (Dietz et al., 2021), and Storybuddy (Dietz et al., 2021; Zhang et al., 2022)).
@@
With increasingly powerful and intelligent capabilities, these AI-based systems and tools have provided tremendous promises for interactive, intelligent, and engaging storytelling experiences for children, acting as their language partners or learning assistants (Xu et al., 2021b; Dietz et al., 2021). Recent booming large-language model-based technologies (LLMs) (Radford et al., 2019) further solidify this vision. This increasing functionality, diversity, and popularity of AI-based technologies in storytelling scenarios for preschoolers raise critical questions regarding: how AI techniques support real-world storytelling practices and how parents, the vital role of storytelling practices, experience, and perceive AI-mediated storytelling for preschoolers?
@@
In CSCW, studies have paid attention to the design and use of various storytelling applications, with the aim of improving children’s digital literacy (Dietz et al., 2021; Tseng et al., 2021), stimulating their creativity and collaboration (Dietz et al., 2021; Helmes et al., 2009), fostering their connection and interaction with parents (Harbig et al., 2011; Zhang et al., 2022), etc. Meanwhile, researchers have evaluated the functionality and effectiveness of existing commercial AI products (e.g., Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021)) as well as the recent LLMs technologies (Han and Cai, 2023b) to support storytelling practices with preschoolers (Xu et al., 2021b; Garg et al., 2022; Chubb et al., 2022). However, these studies were conducted primarily through experimental research. To date, very little has been published on the usage and evaluation of AI-based storytelling technologies in real-world storytelling scenarios beyond a limited research period.
@@
Our study fills this research gap through a qualitative study of 17 parents with preschoolers between the ages of 3 and 6. Our findings indicate that parents have a cautiously positive attitude toward AI-based storytelling. Given the practical challenges facing today’s parents, such as lack of time, energy, and necessary knowledge in carrying out high-quality, meaningful interactive storytelling, our findings suggest that parents expect AI technologies to help alleviate their burdens and support the desired parent-children interactive storytelling. Yet, existing tools generally fail to support parents’ expected storytelling in the real world due to several interactive and algorithmic challenges. Drawing on our findings, we delve into parents’ expectations for ideal parent-child interactive storytelling and propose several crucial design implications for future research.
The key contributions of our work are mainly twofold:
  • We acquire an empirical understanding of parents’ perception and facing challenges regarding the use of AI-based storytelling tools in practice through a qualitative study involving 17 parents with preschoolers.
  • We identify and analyze vital design implications and recommendations to better facilitate the increasing demand for AI-based children-centered storytelling.
@@ 2. Related Work
Our study situates into the research of AI-based storytelling for preschoolers and its practical effectiveness. In this section, we first present a review of existing AI-based storytelling tools. Following that, we review existing literature about the practical usage and evaluation of these tools from children’s and parents’ perspectives respectively.
@@ 2.1. AI-Based Interactive Storytelling Systems
With increasingly powerful and intelligent capabilities, as well as the practical barriers and challenges in traditional parent-child storytelling, there has been a growing CSCW and HCI interest in designing and using AI-based technologies in children’s storytelling scenarios (Frohlich et al., 2009; Dietz et al., 2021; Tseng et al., 2021; Helmes et al., 2009; Harbig et al., 2011; Zhang et al., 2022; Rubegni and Landoni, 2014). Researchers, for instance, have designed storyCoder (Dietz et al., 2021), a voice-guided smartphone application, and PlushPal (Tseng et al., 2021), a web-based tool, to improve children’s computational thinking ability and data literacy by leveraging storytelling as a creative activity. Helmes et al. (2009) designed TellTable, a tabletop storytelling application, to stimulate children’s creativity and collaboration. Meanwhile, researchers have paid specific attention to children with disabilities, such as Gelsomini et al. (2016)’s Wildcard for children with intellectual developmental disabilities and Harbig et al. (2011)’s SignBright for deaf children. Besides, commercially available AI-based robotic toys, such as Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021), etc., also provide platforms for facilitating interactive storytelling experiences for children (Garg and Sengupta, 2020).
@@
The recent booming of Large Language Model technologies (LLMs) further sparked the potential of AI-based storytelling technologies in supporting more intelligent and personalized storytelling practices. Han and Cai (2023b), for instance, examines the design implications of leveraging generative AI tools for improving children’s literacy development and creative story expression. Based on the findings, they proposed AIStory (HAN et al., 2023), a generative AI-powered visual storytelling application prototype that can be used for children’s creative expression, storytelling, and literacy development. Li and Xu (2023) propose a peer-like embodied conversational agent named STARie that integrates multiple AI models that aims to support narrative production in collaborative storytelling, specifically for children aged 4-8.
@@
While these studies provide the worth-expecting vision for AI-mediated storytelling, most existing technologies are still in the prototype design and evaluation phase, involving brief testing for a short time with limited users rather than applying them to real-world scenarios for a long time. Due to the lack of large-scale deployment and long-term usage, it remains unclear how such AI-based storytelling systems are perceived and used by their intended users (e.g. parents and preschoolers) and what barriers and challenges they might encounter. Our study contributes to bridging this gap by examining the use of existing AI-based storytelling tools in the real world, particularly from parents’ perspective.
@@ 2.2. AI-Based Storytelling Technologies and Children
In CSCW and related fields, there has been surging interest in evaluating the performance of AI-based intelligent systems in children’s reading and storytelling scenarios. Researchers, for instance, have illustrated that preschool-aged children can generally interact naturally with conversational agents (CAs), perceiving them as both intelligent and friendly (Druga et al., 2017). Interacting with CAs can also improve children’s story comprehension and learning, comparable to that provided by an adult language partner.(Xu et al., 2021b; Kory and Breazeal, 2014; Beneteau et al., 2020; Xu et al., 2021a; Xu and Warschauer, 2020a). Compared to reading and interacting with human partners, children reading with CAs responded to questions with better intelligibility but lower productivity, lexical diversity, and topical relevance (Xu et al., 2021b).
@@
Meanwhile, literature has also shown that AI applications face difficulties when integrating into real-world storytelling or education practices for preschoolers (Su et al., 2023). From the technical standpoint, researchers, for instance, have demonstrated existing conversational agents are hard to generating the expected valid and reliable comprehension story questions and supporting tailored storytelling and interaction (Xu et al., 2022; Yao et al., 2022; Bentley et al., 2018; Cowan et al., 2017; Luger and Sellen, 2016), which caused children might encounter difficulties in understanding questions from agents, fail to follow the conversation flow, and struggle with appropriate turn-taking when interacting with agents (Du et al., 2021; Pantoja et al., 2019). In addition, many existing AI-based storytelling tools merely ask questions in a sequential manner from a pre-generated list, which does not prioritize child engagement, as it lacks logical coherence between questions and often fails to provide guidance to children in case of confusion or incorrect answers (Zhang et al., 2022; Xu et al., 2022; Yao et al., 2022).
As children are increasingly exposed to various AI technologies, researchers have also explored barriers and limitations of AI technologies to children from a broader perspective of child-AI interaction. They, for instance, examined the potential social and gender biases of AI technologies to children, and proposed to provide them with opportunities and skills to critically reflect on current AI applications and envision (Sharma et al., 2023; Zhou and Yu, 2023). Meanwhile, researchers have also examined the impact of increasing digital technologies and the trend of datafication on children (Wang et al., 2022a, 2023). They suggested future AI technologies should be more transparent and autonomy-supportive to children. Given these children-centered findings, researchers proposed to adapt a generic human-centered AI design framework (Wang et al., 2022b; Zhou and Yu, 2023) for more ethical, inclusive, and fair AI technology designs. These findings and claims could also be reflected in AI storytelling technologies for preschoolers.
@@
Overall, the challenges and limitations identified in prior research raise an urgent need to thoroughly understand users’ usage and experience of AI-based storytelling technologies, and identify the practical challenges and barriers. In particular, preschoolers often lack the essential AI literacy (Su et al., 2023) and skills (Resnick et al., 2009) to independently use AI technologies. These insights are, therefore, more critical in informing the development of future AI-based storytelling technologies for preschoolers, ensuring the alignment with practical storytelling needs, and maximizing benefits in AI-based storytelling. Our work joins the effort of ongoing work to deeply explore the practical challenges and limitations of AI-based storytelling technologies.
@@ 2.3. AI-Based Storytelling Technologies and Parents
For practical storytelling practices with preschoolers, parents play the most vital role in, for instance, shaping children’s language and comprehension skills development, and strengthening the bond between parents and children (Frude and Killick, 2011; Vasalou et al., 2020). In AI-based storytelling practices for preschoolers, parents’ roles and functions are even more significant, because preschoolers often lack the essential literacy (Su et al., 2023) and skills (Resnick et al., 2009) to independently use various AI technologies, and they therefore often rely on parents to navigate this AI-based storytelling process (Zhang et al., 2022). According to the study of Lin et al. (2021), when integrating AI-based intelligent agents into the story reading and telling scenario, about 63% interactions are actually parent-driven, with the majority of parents guiding children by redirecting their attention or clarifying virtual agents’ prompts.
@@
However, researchers have found that existing AI-based storytelling technologies often excluded the involvement of parents (Zhang et al., 2022), and the evaluation of these technologies has also mainly from the children’s needs and perspective, paying attention to children’s cognition of AI technologies and the impact of AI technologies on children’s certain abilities, e.g., (Xu et al., 2021b; Dietz et al., 2021; Cai, 2023a; Gelsomini et al., 2016; Chubb et al., 2022). In general, attention to parents’ experiences and perceptions of AI-based storytelling is limited. In CSCW and HCI, some prior research has provided formative insights into parents’ objectives for interactive storytelling. Zhang et al. (2022), for instance, find that parents appreciate the educational potential of new technologies in enhancing parent-child interaction and actively seeking customized content for their children. Boffi (2020) reports the challenges parent face in educating their children to interact with intelligent storytelling systems correctly and respectfully. Apart from these studies, little prior work has deeply explored parents’ experiences and perceptions of AI-based storytelling technologies.
@@
Given parents’ significant roles in the adopting and using of AI-based storytelling systems (Boffi, 2020; Lin et al., 2021; Zhang et al., 2022), our study provided an empirical study to comprehensively and deeply examine how they perceive and use existing AI-based storytelling tools, their encountered barriers and challenges. Relevant to our study here is Lin et al. (2021)’s study about parental acceptance of children’s storytelling robots. Through a qualitative study with 18 parents, they demonstrate parents’ mixed, though generally positive, attitudes toward children’s storytelling robots. In contrast to their research, our study primarily focused on the challenges and challenges our participants experienced and perceived in real-world settings. Compared to experimental research which occurs in controlled environments, and design fiction (Lindley and Coulton, 2015) which uses fictional scenarios and prototypes to explore and envision possible futures for technology, studying the lived experiences of parents in real-world scenarios can provide us valuable insights into parents’ actual use and perceptions of AI storytelling tools, allowing us to capture the nuances and complexities of how parents integrate such tools into their daily storytelling scenarios, and uncover unforeseen challenges and benefits.
@@
In addition, we paid our attention to more general AI-based storytelling technologies, instead of the specific storytelling robots as what Lin et al. (2021) have done. During our study, we did not specify what kinds of technologies our participants have used, so that the questions could be inclusive to participants to share various levels of experiences with different technologies mediated storytelling practices. Through this broader research scope, we hope to contribute to the general AI-based storytelling technology design by informing designers and developers with in-depth and empirical understandings of the practical challenges and challenges in fitting these technologies into the practical parent-child storytelling scenario and helping them better address these challenges.

### v2

# Exploring Parent’s Needs for Children-Centered AI to Support Preschoolers’ Interactive Storytelling and Reading Activities

@@ Abstract.
Interactive storytelling is vital for preschooler development. While children’s interactive partners have traditionally been their parents and teachers, recent advances in artificial intelligence (AI) have sparked a surge of AI-based storytelling and reading technologies. As these technologies become increasingly ubiquitous in preschoolers’ lives, questions arise regarding how they function in practical storytelling and reading scenarios and, how parents, the most critical stakeholders, experience and perceive these technologies. This paper investigates these questions through a qualitative study with 17 parents of children aged 3-6. Our findings suggest that even though AI-based storytelling and reading technologies provide more immersive and engaging interaction, they still cannot meet parents’ expectations due to a series of interactive and algorithmic challenges. We elaborate on these challenges and discuss the possible implications of future AI-based interactive storytelling technologies for preschoolers.
@@
conference: November 09–13, 2024; San Jose, Costa Rica; copyright: acmlicensedjournal: PACMHCIjournalyear: 2024journalvolume: 8journalnumber: CSCW2article: 496publicationmonth: 11doi: 10.1145/3687035ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Interactive storytelling is a specific form of storytelling where, in addition to merely narrating the story verbatim, parents often sit together with and read to their preschool children111“Preschoolers” or “preschool children” in this paper refers to young children, typically aged 3 to 6, who are in the stage of early childhood education that precedes formal schooling (of Disease Control and Prevention, 2021) about the story content., interact with preschoolers and engage them in guided conversation (Zevenbergen and Whitehurst, 2003) by asking questions and providing responsive feedback. It has been proven to be beneficial to the later language and literacy proficiency of preschoolers, communication skills, cultural and emotional awareness, and other aspects of cognitive development (Wright, 1995; Yen et al., 2018; Chang and Huang, 2016; Peck, 1989). However, due to various practical challenges, today’s parents may not always have sufficient language skills, knowledge, time, or inclination to engage in such conversation-rich storybook reading with their preschoolers, even though they believe in the importance of interactive storytelling with their children (Manz et al., 2010; Xu et al., 2021b).
@@
In recent years, AI-based tools and systems with multiple functions, modalities, and interaction techniques have become increasingly popular in supporting interactive storytelling practices for preschoolers. These tools can be applications on mobile phones or tablets (Dietz et al., 2021; Zhang et al., 2022; Zhao et al., 2022), or exist with physical anthropomorphic tangible presences (e.g., Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021)), relying on interactive technologies tailored to children (e.g., (Xu et al., 2022; Yao et al., 2022)), or general voice assistant technologies (e.g., Amazon Alexa, Google Assistant, Apple Siri). Some of these tools are mature commercial products (e.g., Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021)) and some others are still in the research period (e.g., StoryCoder (Dietz et al., 2021), and StoryBuddy (Dietz et al., 2021; Zhang et al., 2022)). With increasingly powerful and intelligent capabilities, these AI-based systems and tools have provided tremendous promises for interactive, intelligent, and engaging storytelling and reading experiences for children, acting as their language partners or learning assistants (Xu et al., 2021b; Dietz et al., 2021). Recent boom in large language model (LLM)-based technologies (Radford et al., 2019) further solidify this vision (Mahmood et al., 2023). This increasing functionality, diversity, and popularity of AI-based technologies in interactive storytelling scenarios for preschoolers raise critical questions regarding: how AI techniques support real-world interactive storytelling and reading practices and how parents, the vital role of storytelling practices, experience, and perceive AI-mediated storytelling for preschoolers?
@@
In CSCW, studies have paid attention to the design, development, and use of various storytelling applications, with the aim of improving children’s digital literacy (Dietz et al., 2021; Tseng et al., 2021), stimulating their creativity and collaboration (Dietz et al., 2021; Helmes et al., 2009), fostering their connection and interaction with parents (Harbig et al., 2011; Zhang et al., 2022), etc. Meanwhile, researchers have evaluated the functionality and effectiveness of existing commercial AI products (e.g., Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021)) as well as the recent LLM technologies (Han and Cai, 2023b; Chen et al., 2023) in supporting storytelling practices with preschoolers (Xu et al., 2021b; Garg et al., 2022; Chubb et al., 2022). However, these studies were conducted primarily through experimental research. To date, very little has been published on the usage and evaluation of AI-based storytelling and reading technologies in real-world interactive storytelling scenarios beyond a limited research period.
@@
Our study fills this research gap through a qualitative study of 17 parents with preschoolers between the ages of 3 and 6. Our findings indicate that parents in our study have a cautiously positive attitude towards AI-based interactive storytelling technologies. Given the practical challenges facing today’s parents in carrying out high-quality and meaningful interactive storytelling, our findings suggest that parents expect AI technologies to help alleviate their burdens and support the desired parent-children interactive storytelling. Yet, existing tools generally fail to support parents’ expected storytelling in the real world due to several interactive and algorithmic challenges. Drawing on our findings, we delve into parents’ expectations for ideal parent-child interactive storytelling and propose several crucial design implications for future research. Our study contributes to the CSCW community by 1) acquiring an empirical understanding of parents’ perception and facing challenges regarding the use of AI-based storytelling tools in practice through a qualitative study involving 17 parents with preschoolers, and 2) identifying and analyzing vital design implications and recommendations to better facilitate the increasing demand for AI-based children-centered storytelling.
@@ 2. Related Work
Our study situates itself in the research of AI-based storytelling for preschoolers and its practical effectiveness. In this section, we first present a review of existing AI-based storytelling tools. Following that, we review existing literature about the practical usage and evaluation of these tools from children’s and parents’ perspectives, respectively.
@@ 2.1. AI-Based Interactive Storytelling Systems
With increasingly powerful and intelligent capabilities, as well as the practical barriers and challenges in traditional parent-child storytelling, there has been a growing interest in CSCW and HCI in designing and using AI-based technologies in children’s storytelling scenarios (Frohlich et al., 2009; Dietz et al., 2021; Tseng et al., 2021; Helmes et al., 2009; Harbig et al., 2011; Zhang et al., 2022; Rubegni and Landoni, 2014). Researchers, for instance, have designed StoryCoder (Dietz et al., 2021), a voice-guided smartphone application, and PlushPal (Tseng et al., 2021), a web-based tool, to improve children’s computational thinking ability and data literacy by leveraging storytelling as a creative activity. Helmes et al. (2009) designed TellTable, a tabletop storytelling application, to stimulate children’s creativity and collaboration. Harbig et al. (2011) and Zhang et al. (2022) designed SignBright and StoryBuddy, respectively, with the aim of fostering connection, collaboration, and understanding between children and parents in storytelling scenarios. Cai (2023a) designed EmotionBlock, a tangible toolkit for children-oriented social-emotional learning through storytelling. Meanwhile, researchers have paid specific attention to children with disabilities, such as Gelsomini et al. (2016)’s Wildcard for children with intellectual developmental disabilities and Harbig et al. (2011)’s SignBright for deaf children. Besides, commercially available AI-based robotic toys, such as Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021), etc., also provide platforms for facilitating interactive storytelling experiences for children (Garg and Sengupta, 2020).
@@
The recent boom of Large Language Model (LLM) technologies further sparked the potential of AI-based technologies, particularly with a verbal communication interface (Mahmood et al., 2023; Yang et al., 2024a; Wan et al., 2024; Wu et al., 2024), in supporting more intelligent and personalized storytelling practices. Han and Cai (2023b), for instance, examines the design implications of leveraging generative AI tools to improve children’s literacy development and creative story expression. Based on the findings, they proposed AIStory (HAN et al., 2023), a prototype of a generative AI-powered visual storytelling application that can be used for children’s creative expression, storytelling, and literacy development. Li and Xu (2023) propose a peer-like embodied conversational agent named STARie that integrates multiple AI models that aims to support narrative production in collaborative storytelling, specifically for children aged 4-8.
@@
While these studies provide the worth-expecting vision for AI-mediated storytelling, most existing technologies are still in the prototype design and evaluation phase, involving brief testing for a short time with limited users rather than applying them to real-world scenarios for a long time. Due to the lack of large-scale deployment and long-term usage, it remains unclear how such AI-based storytelling systems are perceived and used by their intended users (e.g., parents and preschoolers) and what barriers and challenges they might encounter (e.g., ethical concerns (Isaza et al., 2023) or trust in AI issues (Drozdal et al., 2020)). In light of the trend of AI-mediated storytelling for preschoolers, urgent attention is needed to deeply understand these aspects, which can help the designers and developers of AI-based storytelling systems better design these technologies integrate into the practical storytelling scenario to form a “human-parent collaboration” team (Wang et al., 2019, 2020). Our study contributes to bridging this gap by examining the use of existing AI-based storytelling tools in the real world, particularly from parents’ perspective.
@@ 2.2. AI-Based Storytelling Technologies and Children
In CSCW and related fields, there has been a surging interest in evaluating the performance of AI-based intelligent systems in children’s reading and storytelling scenarios. Researchers, for instance, have illustrated that preschool-age children can generally interact naturally with conversational agents (CAs), perceiving them as intelligent and friendly (Druga et al., 2017). Interacting with CAs can also improve children’s story comprehension and learning (Xu et al., 2021b; Kory and Breazeal, 2014; Beneteau et al., 2020; Xu et al., 2021a; Xu and Warschauer, 2020a). Compared to reading and interacting with human partners, children reading with CAs respond to questions with better intelligibility but lower productivity, lexical diversity, and topical relevance (Xu et al., 2021b).
@@
Despite consistent findings that demonstrate the potential and benefits of AI technologies to support storytelling practices for children, literature has also shown that AI applications face difficulties in integrating into real-world storytelling or education practices for preschoolers (Su et al., 2023). From the technical standpoint, researchers, for instance, have demonstrated existing conversational agents face challenges in generating the expected valid and reliable comprehension story questions and supporting tailored storytelling and interaction (Xu et al., 2022; Yao et al., 2022; Bentley et al., 2018; Cowan et al., 2017; Luger and Sellen, 2016; Yang et al., 2024b), which caused children to encounter difficulties in understanding questions from agents, fail to follow the conversation flow, and struggle with appropriate turn-taking when interacting with agents (Du et al., 2021; Pantoja et al., 2019). In addition, many existing AI-based storytelling tools merely ask questions in a sequential manner from a pre-generated list, which does not prioritize child engagement, as it lacks logical coherence between questions and often fails to provide guidance to children in case of confusion or incorrect answers (Zhang et al., 2022; Xu et al., 2022; Yao et al., 2022). In addition, researchers have also pointed out that existing technologies often exclude the participation of parents (Zhang et al., 2022), who actually have significant roles in practical storytelling (Frude and Killick, 2011; Vasalou et al., 2020).
@@
The challenges and limitations identified in prior research raise an urgent need to thoroughly understand users’ usage and experience of AI-based storytelling technologies, and identify the practical challenges and barriers. In particular, preschoolers often lack the essential AI literacy (Su et al., 2023) and skills (Resnick et al., 2009) to use AI technologies independently. These insights are, therefore, more critical to inform the development of future AI-based storytelling technologies for preschoolers, ensuring the alignment with practical storytelling needs and maximizing benefits in AI-based storytelling. Our work joins the ongoing effort to explore the practical challenges and limitations of AI-based storytelling technologies.
Further, as children are increasingly exposed to various AI technologies, researchers have also explored barriers and limitations of AI technologies to children from a broader perspective of child-AI interaction (UNICEF, 2021). Researchers examined the potential social and gender biases of AI technologies to children, and proposed to provide them with opportunities and skills to critically reflect on current AI applications and envision (Sharma et al., 2023; Zhou and Yu, 2023; Wang et al., 2024). Meanwhile, researchers have also examined the impact of increasing digital technologies and the trend of datafication in children (Wang et al., 2022a, 2023). They suggested that future AI technologies should be more transparent and autonomy-supportive to children. Given these children-centered findings, researchers proposed to adapt a generic human-centered AI design framework (Wang et al., 2022b; Zhou and Yu, 2023) for more ethical, inclusive, and fair AI technology designs. These findings and claims could also be reflected in AI storytelling technologies for preschoolers. Drawing from the notion of child-centered AI design (UNICEF, 2021), our study deeply examines the usage and impacts of AI-based interactive storytelling technologies, with a particular focus on human (i.e., parents and children) needs, experiences, and perceptions. According to our findings, we proposed a series of human-centered design implications for future human-centered, AI-based interactive storytelling technologies.
@@ 2.3. AI-Based Storytelling Technologies and Parents
For practical storytelling practices with preschoolers, parents play the most vital role in, for instance, shaping children’s language and comprehension skills development, and strengthening the bond between parents and children (Frude and Killick, 2011; Vasalou et al., 2020). In AI-based storytelling practices for preschoolers, parents’ roles and functions are even more significant, because preschoolers often lack the essential literacy (Su et al., 2023) and skills (Resnick et al., 2009) to independently use various AI technologies, and they therefore often rely on parents to navigate this AI-based storytelling process (Zhang et al., 2022). According to the study of Lin et al. (2021), when integrating AI-based intelligent agents into the story reading and telling scenario, about 63% interactions are actually driven by parents, and the majority of parents guide children by redirecting their attention or clarifying virtual agents’ prompts.
@@
However, existing AI-based storytelling technologies have been mainly developed and evaluated from children’s needs and perspectives (Xu et al., 2021b; Dietz et al., 2021; Cai, 2023a; Gelsomini et al., 2016; Chubb et al., 2022). In general, attention to parents’ experiences and perceptions of AI-based storytelling is limited. In CSCW and HCI, some prior research has provided formative insight into parents’ objectives for interactive storytelling. Zhang et al. (2022), for instance, find that parents appreciate the educational potential of new technologies in enhancing parent-child interaction and actively seeking customized content for their children. Boffi (2020) reports the challenges parents face in educating their children to interact with intelligent storytelling systems correctly and respectfully. Apart from these studies, little prior work has deeply explored parents’ experiences and perceptions of AI-based storytelling technologies.
@@
Given parents’ significant roles in the adopting and using of AI-based storytelling systems (Boffi, 2020; Lin et al., 2021; Zhang et al., 2022), our study provided an empirical study to comprehensively and deeply examine how they perceive and use existing AI-based storytelling tools, their encountered barriers and challenges. Relevant to our study here is Lin et al. (2021)’s study about parental acceptance of children’s storytelling robots. Through a qualitative study with 18 parents, they demonstrate mixed, though generally positive attitudes of parents towards children’s storytelling robots. In contrast to their research, our study primarily focused on the challenges our participants experienced and perceived in real-world settings. Compared to experimental research that occurs in controlled environments and design fiction (Lindley and Coulton, 2015) which uses fictional scenarios and prototypes to explore and envision possible futures for technology, studying the lived experiences of parents in real-world scenarios can provide us with valuable insights into parents’ actual use and perceptions of AI storytelling tools, allowing us to capture the nuances and complexities of how parents integrate such tools into their daily storytelling scenarios, and uncover unforeseen challenges and benefits. These in-depth, practical insights can help the designers and developers of AI-based storytelling tools to better align with the practical needs and preferences of parents, and understand and address potential issues in fitting AI-based storytelling tools into their storytelling practices.
@@
In addition, we paid our attention to more general AI-based storytelling technologies, instead of the specific storytelling robots like what Lin et al. (2021) have done. During our study, we did not specify what kinds of technologies our participants have used, so that the questions could be inclusive of participants to share various levels of experience with different technologies-mediated storytelling practices. Through this broader research scope, we hope to contribute to the general AI-based storytelling technology design by informing designers and developers with in-depth and empirical understandings of the practical challenges and challenges in fitting these technologies into the practical parent-child storytelling scenario and helping them better address these challenges.